# Graph Generative Pre-trained Transformer

### v1

@@ 5.1 Setup
Datasets. We consider both generative tasks and predictive tasks in our experiments. In generative tasks, we consider training transformer decoders on molecular datasets and generic graph datasets. For molecular datasets, we use QM9 (Wu et al., 2018b), MOSES (Polykovskiy et al., 2020), and GuacaMol (Brown et al., 2019). For generic graph datasets, we adapted the widely used datasets: Planar, Tree, Lobster, and stochastic block model (SBM). In predictive tasks, we fine-tune models pre-trained from GuacaMol datasets on various molecular property tasks using MoleculeNet (Wu et al., 2018a), detailed in Section B.4, to verify the usefulness of the learned graph representations.
Model specifications. We train transformers with three different sizes: (1) the small transformer has 6 transformer layers and 6 attention head, with , leading to approximately 10M parameters; (2) the base transformer has 12 transformer layers and 12 attention head, with , leading to approximately 85M parameters; (2) the large transformer has 24 transformer layers and 16 attention head, with , leading to approximately 300M parameters. We use different specifications for different experiments according to the task complexity.
5.2 A Demonstrative Experiment using Planar Graphs
We first validate the effectiveness of our proposed graph sequence representation compared to the adjacency matrix. To achieve this, we train transformer decoders on planar graphs using both representations and evaluate their generative performance. For the adjacency representation, planar graphs are encoded as sequences of 0s and 1s derived from the strictly lower triangular matrix, with rows and columns permuted using BFS orderings to augment the training dataset. Table 3 presents the quantitative and qualitative results of the generated samples. Our proposed representation demonstrates superior generative performance with a much smaller set of tokens, while model learning adjacency matrices struggles to capture the topological rules of the training graphs.
@@ 5.7 Scaling Effects
For model scaling, we additionally train G2PTs with 1M, 705M, and 1.5B parameters. We notice that as model size increases, validity score generally increases and saturates at some point, depending on the task complexity. For instance, QM9 saturates at the beginning (1M parameters) while MOSES and GuacaMol require more than 85M (base) parameters to achieve satisfying performance.
@@ 6 Conclusion
This work revisits the graph generative models and proposes a novel sequence-based representation that efficiently encodes graph structures via node and edge definitions. This representation serves as the foundation for the proposed Graph Generative Pre-trained Transformer (G2PT), an auto-regressive model that effectively models graph sequences using next-token prediction. Extensive evaluations demonstrated that G2PT achieves remarkable performance across multiple datasets and tasks, including generic graph and molecule generation, as well as downstream tasks like goal-oriented graph generation and graph property prediction. The results highlight G2PT’s adaptability and scalability, making it a versatile framework for various applications. One limitation of our method is that G2PT is order-sensitive, where different graph domains may prefer different edge orderings. Future research could be done by exploring edge orderings that are more universal and expressive.
@@ B.5 Baselines
Graph pre-training methods. We compare against several pre-training approaches for molecular property prediction, as summarized in Table 5. The goal ofGraph pre-training methods is to learn robust graph representations via exploiting the structural information. AttrMask (Hu et al., 2020) uses attribute masking at both node and graph levels to capture local and global features simultaneously. ContextPred (Hu et al., 2020) builds on this idea by predicting subgraph contexts, enabling the model to understand patterns beyond individual attributes. Similarly, InfoGraph (Sun et al., 2020) focuses on multi-scale graph representations by maximizing mutual information between graph-level embeddings and substructures. Moving to contrastive learning approaches, GraphCL (You et al., 2021) applies graph augmentations to generate positive and negative samples for representation learning. Building on this idea, GraphMVP (Liu et al., 2022a) incorporates both 2D molecular topology and 3D geometric views, aligning them within a contrastive framework to enhance feature representation. In contrast to these methods, GraphMAE (Hou et al., 2022b) adopts a generative approach, using a masked graph auto-encoder to reconstruct node features and capture structural information.
@@ B.6 Evaluation
Appendix C Related works
@@ C.2 Non-auto-regressive Graph Generative Models
In addition to auto-regressive models, non-auto-regressive graph generative models can be categorized into two branches: (1) one-shot generation and (2) iterative refinement. One-shot generation aims to generate a graph in a single step including methods such as generative adversarial networks (De Cao & Kipf, 2018), variational auto-encoders (Simonovsky & Komodakis, 2018; Liu et al., 2018) and normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020). Nevertheless, one-shot graph generative models often suffer from the decoding strategies such that it requires an expressive decoder to map from latent vectors to graphs. On the other side, iterative refinement methods generate the entire graph in the first step and then iteratively refine the generated graph to be close to a realistic graph, including diffusion (Niu et al., 2020; Jo et al., 2022a; Vignac et al., 2022; Chen et al., 2022, 2023; Jo et al., 2023; Haefeli et al., 2022; Wu et al., 2023; Siraudin et al., 2024; Xu et al., 2024) and flow matching models (Qin et al., 2024; Eijkelboom et al., 2024; Lipman et al., 2022; Liu et al., 2022b; Esser et al., 2024; Ma et al., 2024; Campbell et al., 2024; Gat et al., 2024). As discussed in Section 2, they often require a prefixed number of refinement steps and they need to maintain an adjacency matrix over the trajectory which is computationally intensive.
@@ D.2 Additional Visualizations
![[Uncaptioned image]](extracted/6106521/images/demo_planar_A_1.png)
![[Uncaptioned image]](extracted/6106521/images/demo_planar_A_2.png)
![[Uncaptioned image]](extracted/6106521/images/demo_planar_ours_1.png)
![[Uncaptioned image]](extracted/6106521/images/demo_planar_ours_2.png)
![[Uncaptioned image]](extracted/6106521/images/moses_train_1.png)
![[Uncaptioned image]](extracted/6106521/images/moses_10m_1.png)
![[Uncaptioned image]](extracted/6106521/images/guac_train_1.png)
![[Uncaptioned image]](extracted/6106521/images/guac_10m_1.png)
![[Uncaptioned image]](extracted/6106521/images/moses_train_2.png)
![[Uncaptioned image]](extracted/6106521/images/moses_10m_2.png)
![[Uncaptioned image]](extracted/6106521/images/guac_train_2.png)
![[Uncaptioned image]](extracted/6106521/images/guac_10m_2.png)

### v2

@@ 5.1 Setup
Datasets. We consider both generative tasks and predictive tasks in our experiments. We use three molecule datasets: QM9 (Wu et al., 2018b), MOSES (Polykovskiy et al., 2020), and GuacaMol (Brown et al., 2019); and four generic graph datasets: Planar, Tree, Lobster, and stochastic block model (SBM), which are widely used to benchmark graph generative models. In predictive tasks, we fine-tune models pre-trained from GuacaMol datasets on molecular properties with the benchmark method MoleculeNet (Wu et al., 2018a). Further details are in Section B.4,
Model specifications. We train Transformers with three different sizes: (1) the small Transformer has 6 layers and 6 attention heads, with , leading to approximately 10M parameters; (2) the base Transformer has 12 layers and 12 attention heads, with , leading to approximately 85M parameters; (2) the large Transformer has 24 layers and 16 attention heads, with , leading to approximately 300M parameters. We use different specifications for different experiments according to the task complexity.
5.2 A Case Study with Planar Graphs
We compare our token-based representation against the adjacency matrix and validate its effectiveness in the generation task. We train Transformer decoders on planar graphs using our token-based representation and adjacency matrices, and then evaluate their generative performance. For the adjacency representation, planar graphs are encoded as sequences of 0s and 1s derived from the strictly lower triangular matrix, with rows and columns permuted using BFS orderings to augment the training dataset. Table 3 presents the quantitative and qualitative results of the generated samples. Our proposed representation demonstrates superior generative performance with a much smaller set of tokens. In contrast, the model trained with adjacency matrices struggles to capture the topological rule of planar graphs.
@@ 5.7 Scaling Effects
For model scaling, we additionally train G2PTs with 1M, 707M, and 1.5B parameters. We notice that as model size increases, validity score generally increases and saturates at some point, depending on the task complexity. For instance, QM9 saturates at the beginning (1M parameters) while MOSES and GuacaMol require more than 85M (base) parameters to achieve satisfying performance.
@@ 6 Conclusion
This work revisits the sequential approach to graph generation and proposes a novel token-based representation that efficiently encodes graph structures via node and edge tokens. This representation serves as the foundation for the proposed Graph Generative Pre-trained Transformer (G2PT), an auto-regressive model that effectively models graph sequences using next-token prediction. Extensive evaluations demonstrated that G2PT achieves remarkable performance across multiple datasets and tasks, including generic graph and molecule generation, as well as downstream tasks like goal-oriented graph generation and graph property prediction. The results highlight G2PT’s adaptability and scalability, making it a versatile framework for various applications. One limitation of our method is that G2PT is order-sensitive, where different graph domains may prefer different edge orderings. Future research could be done by exploring edge orderings that are more universal and expressive.
Impact Statement
This paper introduces a framework that models graphs in a similar vein to GPT (Generative Pre-trained Transformer). The G2PT framework allows seamless plantation of training techniques that have developed in other domains based on GPT. Besides performing generative tasks such as drug discovery, G2PT also can be easily extended for discriminative tasks such as graph property prediction. We hope this work will advance the field of graph learning. As a powerful tool, G2PT may also be used as one step in a complex system to create molecule structures harmful to humans or the environment, but we don’t see immediate hazards from our study.
Acknowledgment
We thank all reviewers for their insightful feedback. Chen and Liu’s work was supported by NSF 2239869. He and Xu’s work was support by NSF 2239672.
@@ B.5 Baselines
Graph pre-training methods. We compare against several pre-training approaches for molecular property prediction, as summarized in Table 5. The goal ofGraph pre-training methods is to learn robust graph representations via exploiting the structural information. AttrMask (Hu et al., 2020a) uses attribute masking at both node and graph levels to capture local and global features simultaneously. ContextPred (Hu et al., 2020a) builds on this idea by predicting subgraph contexts, enabling the model to understand patterns beyond individual attributes. Similarly, InfoGraph (Sun et al., 2020) focuses on multi-scale graph representations by maximizing mutual information between graph-level embeddings and substructures. Moving to contrastive learning approaches, GraphCL (You et al., 2021) applies graph augmentations to generate positive and negative samples for representation learning. Building on this idea, GraphMVP (Liu et al., 2022a) incorporates both 2D molecular topology and 3D geometric views, aligning them within a contrastive framework to enhance feature representation. In contrast to these methods, GraphMAE (Hou et al., 2022b) adopts a generative approach, using a masked graph auto-encoder to reconstruct node features and capture structural information.
@@ B.6 Evaluation
Appendix C Extended Related works
@@ C.2 Non-auto-regressive Graph Generative Models
In addition to auto-regressive models, non-auto-regressive graph generative models can be categorized into two branches: (1) one-shot generation and (2) iterative refinement. One-shot generation aims to generate a graph in a single step including methods such as generative adversarial networks (De Cao & Kipf, 2018), variational auto-encoders (Simonovsky & Komodakis, 2018; Liu et al., 2018) and normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020). Nevertheless, one-shot graph generative models often suffer from the decoding strategies such that it requires an expressive decoder to map from latent vectors to graphs. On the other side, iterative refinement methods generate the entire graph in the first step and then iteratively refine the generated graph to be close to a realistic graph, including diffusion (Niu et al., 2020; Jo et al., 2022a; Vignac et al., 2022; Chen et al., 2022b, 2023; Jo et al., 2023; Haefeli et al., 2022; Wu et al., 2023; Siraudin et al., 2024; Xu et al., 2024; Wang et al., 2025) and flow matching models (Qin et al., 2024; Eijkelboom et al., 2024; Lipman et al., 2022; Liu et al., 2022b; Campbell et al., 2024; Gat et al., 2024). As discussed in LABEL:sec:preli, they often require a prefixed number of refinement steps and they need to maintain an adjacency matrix over the trajectory which is computationally intensive.
C.3 Pre-training Transformers for Graphs
Transformers are now dominating domains of natural language processing (NLP) and computer vision (CV) (Radford, 2018; Devlin, 2018; Dosovitskiy, 2020). Several works also attempt to applying transformers in the field of graph learning (Ying et al., 2021; Hu et al., 2020b; Dwivedi & Bresson, 2020; Rampášek et al., 2022; Chen et al., 2022a; Wu et al., 2021; Kreuzer et al., 2021; Min et al., 2022; Chen et al., 2024; Gao et al., 2024). Those approaches propose several methods to encode the graph structure information into sequences, specifically, the key research problem lies in how to add identifiers to nodes and tokenize the edges. For instance, Kim et al. (2022) uses positional embedding to help transformer to identify nodes, and type embeddings to distinguish node and edge tokens. A more recent work, Gao et al. (2024), which focuses on molecule representation learning, uses a vocabulary to store all atom types and all possible bond types (same bonds with different atoms as endpoints are considered as different type in their case). In contrast, by introducing node index into the vocabulary, G2PT easily implements the edge tokenizations and node identifications.
@@ D.2 Additional Visualizations
![[Uncaptioned image]](extracted/6509401/images/demo_planar_A_1.png)
![[Uncaptioned image]](extracted/6509401/images/demo_planar_A_2.png)
![[Uncaptioned image]](extracted/6509401/images/demo_planar_ours_1.png)
![[Uncaptioned image]](extracted/6509401/images/demo_planar_ours_2.png)
![[Uncaptioned image]](extracted/6509401/images/moses_train_1.png)
![[Uncaptioned image]](extracted/6509401/images/moses_10m_1.png)
![[Uncaptioned image]](extracted/6509401/images/guac_train_1.png)
![[Uncaptioned image]](extracted/6509401/images/guac_10m_1.png)
![[Uncaptioned image]](extracted/6509401/images/moses_train_2.png)
![[Uncaptioned image]](extracted/6509401/images/moses_10m_2.png)
![[Uncaptioned image]](extracted/6509401/images/guac_train_2.png)
![[Uncaptioned image]](extracted/6509401/images/guac_10m_2.png)